OpenStudy (thehouseisaround):

Find the radius, diameter and circumference of a circle with an area of 36pi symbol cm2. Write your answers as decimals rounded to the nearest hundredth.

TheSmartOne (thesmartone):

\(area = \pi r^2\) \(area = 36\pi\) \(36\pi = \pi r^2\) solve for r, which is radius :)

OpenStudy (thehouseisaround):

can you please explain how to do it because I really have no clue about this stuff thank you

OpenStudy (triciaal):